OpenStudy (anonymous):

Hey would i wrote the equation for this p,ease explain in second grader terms=)

OpenStudy (venomblast):

where the question?

OpenStudy (anonymous):

how would i write an equatio n for that...im very confused..

OpenStudy (venomblast):

so that a negative slope

OpenStudy (anonymous):

y=mx+b m=slope=rise over run= y/x

OpenStudy (anonymous):

plug in x*slope and minus it from y to get b and there is your equation

OpenStudy (anonymous):

uuuuuh? y=mx+b this doesnt make much sense...all the possible answers have like a greater than or equal to sign or a less than or equal to sign or < or >

OpenStudy (anonymous):

?

OpenStudy (nubeer):

the region below the line is shaded.. that means your equation will be less then ... and the line drawn is solid that means less then equal to..

OpenStudy (nubeer):

in case that line drawn would be dottted then you would just put less then sign.

OpenStudy (anonymous):

slope= -1/2 y intercept = 4 y=-1/2x+4

OpenStudy (anonymous):

like whats the blue print for the equation i have to write like y=mx+b but like only in that iother for

OpenStudy (anonymous):

it's an inequality, 1 second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

instead of y = it will be < or > or equal to or greater than/ less than that all depends on the shaded part of the graph

OpenStudy (anonymous):

like nubeer said shaded below the line less than and line is solid equal too.

OpenStudy (anonymous):

uuuh im so sorry im just very confused can you start over and make it a little more simple in steps i promise to fan and give you a medal if you do!

OpenStudy (anonymous):

\[y \le -\frac{ 1 }{ 2 } x + 4\]